A coffee merchant has coffee beans that sells for $9 per pound and $12 per pound. The two types are to be mixed to create 100 lb. of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
Answer by venugopalramana(3286)

A coffee merchant has coffee beans that sells for $9 per pound and $12 per pound. The two types are to be mixed to create 100 lb. of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
LET QTY OF $9 TYPE IN MIX = N LB.
QTY OF $ 12 TYPE IN MIX = 100-N LB....SINCE TOTAL MIX IS 100LB.
CHARGE FOR $ 9 TYPE = 9N
CHARGE FOR $ 12 TYPE = 12(100-N)
TOTAL CHARGE = 9N+12(100-N)= 9N+1200-12N = 1200-3N
THIS SHOULD COST $ 11.25 / LB..OR...100*11.25=$ 1125
1125 = 1200-3N
3N = 1200-1125=75
N=25
HENCE 25 LBS OF $ 9 TYPE AND 75 LBS OF $ 12 TYPE ARE TO BE BLENDED.